The financial records of LeRoi Jones Inc. were destroyed by fire at the end of 2020. Fortunately, the controller had kept certain statistical data related to the income statement as follows.

1. The beginning merchandise inventory was $92,000 and decreased 20% during the current year.
2. Sales discounts amount to $17,000.
3. 20,000 shares of common stock were outstanding for the entire year.
4. Interest expense was $20,000.
5. The income tax rate is 30%.
6. Cost of goods sold amounts to $500,000.
7. Administrative expenses are 20% of cost of goods sold but only 8% of gross sales.
8. Four-fifths of the operating expenses relate to sales activities.

From the foregoing information prepare an income statement for the year 2020 in single-step form.
Answer #1
Income Statement
Particulars Amount Amount
Sales $1,250,000.00
Less:Sales Discount -$17,000.00
Net Sales $1,233,000.00
Less: Expenses
Cost Of Goods Sold $500,000.00
Administrative Expenses $100,000.00
Selling Expenses $400,000.00
Interest Expenses $20,000.00
Total Expenses $1,020,000.00
Income Before Taxes $213,000.00
Taxes @30% $63,900.00
Net income $149,100.00
Earnings Per Share (149100/20000) $7.46
Administrative Expenses= 20% X Cost of Goods sold
=500000*20%= $100000
Selling Expenses = Administrative Expenses x 4
=$100,000 x 4= $400000
Gross sales= Administrative Expenses/8%
=100000/8%=1250000